Status Report

ISS On-Orbit Status 15 Jan 2003

By SpaceRef Editor
January 15, 2003
Filed under , ,
ISS On-Orbit Status 15 Jan 2003

All ISS systems continue to function nominally, except as noted previously
or below. A highly successful "golden-50th" EVA day.

Shortly after crew wake-up (1:00am, all times EST), MCC-H began preparations
for the translation of the MT (mobile transporter) trolley from its current
WS7 (work station 7) at the end of the P1 truss to its WS4 home station
on the S0. The move of the rail car got underway at 2:55am and was successfully
completed at 3:10am with latches in place. During the activities, structural
dynamics data were taken with the SDMS (structural dynamics measurement
system) on the S0 truss.

Meanwhile, at 2:05am, after breakfast, the crew started EVA preparations
and ingressed the Airlock for 10.2-psi operations. Assisted by FE-1 Nikolai
Budarin, CDR Ken Bowersox and FE-2/SO Don Pettit donned EMUs and purged
the suits, after which Budarin repressurized the A/L to 14.7 psi and
opened the A/L-Node hatch to allow the spacewalkers to conduct a one-hour
O2 pre-breathe while exercising. The A/L crew lock was then depressurized
to vacuum. Hatch opening by Sox was delayed by 20 minutes due to interference
by the thermal hatch cover fabric.

The Stage EVA began at 7:50am when the EMUs were switched to internal
power, and ended at 2:41pm, after a duration of 6h 51m, with almost all
objectives completed. [With Budarin acting as IV (intravehicular crewmember),
Bowersox and Pettit conducted all tasks smoothly and without hitch, except
for the installation of a light fixture on a boom stanchion on one of
the two CETA (crew equipment translation aid) railcarts on S1, due to
a problem with boom rotation (a pin not dislodging). This low-priority
task was deferred to a future spacewalk. Sox and Don completed the release
of 10 launch restraint locks, monitored the successful deployment and
lock-down of the huge central P1 radiator panel (75 ft), cleaned off
the Node nadir ACBM (active common berthing module) docking surface with
sticky tape, conducted a systems check of the filled ammonia coolant
reservoir on the P6 truss, and relocated & inventoried Z1 and A/L
tool box tools. After EVA cleanup, the crew returned to the A/L, cut
off a fabric loop on the inside of the thermal hatch cover that had caused
the initial hang-up, and completed the spacewalk. It was the 50th EVA
of station assembly, the 25th from the ISS (the others from the Shuttle),
and the 16th from the Joint A/L (the others from the Russian DC-1), but
only the second with no Shuttle present. Total US EVA time for ISS is
now 312h 11m (144h 56m from the A/L).]

After Budarin had assisted the spacewalkers with suit doffing and other
post-EVA activities, he prepared the "Urolux" equipment for
tomorrow’s scheduled post-EVA session of the Russian biochemical urine
test (PZE MO-9).

Earlier in the day, Bowersox had performed the daily Lab payloads status
checkup, while Budarin completed his daily inspection of the Russian
BIO-5 Rasteniya-2/Lada-2 ("Plants-2") experiment.

The MCA (major constituents analyzer) has been working again since yesterday
and was calibrated successfully, alleviating the need for daily manual
O2 and CO2 data takes by the crew.

Crew sleep period begins at the regular 4:30pm EST.

Later tonight, at 6:00pm, attitude control authority will be handed
over to Russian MCS (motion control system) thrusters for maneuvering
the station from the current LVLH (local vertical/local horizontal) "earth-fixed" attitude
to solar-oriented XPOP (x-axis perpendicular to orbit plane) at 6:15pm.
Half an hour later, attitude control will be returned to U.S. CMG momentum

Today’s (optional) targets for the CEO (crew earth observations program)
were Dakar, Senegal (looking right for Dakar; and left for a well-formed,
feathered jet stream cloud, stretching from Senegal to Egypt [related
dynamically to the present storm in the Mediterranean]), St. Thomas Island,
Caribbean (nadir pass. Of interest: coral reef detail), St. John Island,
Caribbean(nadir pass. Of interest: coral reef detail), St. Croix Island,
Caribbean (nadir pass. Of interest: coral reef detail) and Lower Amazon
River Basin (views left and right of the enormous estuary of the planet’s
largest river. Locations and shapes of islands in the estuary are now
known to change all the time).

CEO images can be viewed at the website

U.S. and Russian Segment Status (as of 1:15pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):
Elektron O2 generator is powered On (16 Amp mode). Vozdukh CO2 scrubber
is Off (but functional). U.S. CDRA CO2 scrubber is Off. TCCS (trace
contaminant control subsystem) is operational. MCA (major constituents
analyzer) is operating. BMP Harmful Impurities unit: Absorbent bed
#1 in Purify mode, bed #2 in Purify mode.

SM Working Compartment: Pressure (mmHg) — 753; temperature (deg C)
— 22.3; ppO2 (mmHg) — 159.3; ppCO2 (mmHg) — 2.4 (suspect).
SM Transfer Compartment: Pressure (mmHg) — 751; temperature (deg C)
— 18.6.
FGB Cabin: Pressure (mmHg) — 752; temperature (deg C) — 22.0.
Node: Pressure (mmHg) — 748.98; temperature (deg C) — 23.1 (shell);
ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
U.S. Lab: Pressure (mmHg) — 751.18; temperature (deg C) — 22.6; ppO2
(mmHg) — n/a; ppCO2 (mmHg) — n/a;
Joint Airlock (Equip. Lock): Pressure (mmHg) — 751.18, temperature (deg
C) — 21.1; shell heater temp (deg C) — 21.6, ppO2 (mmHg) — 164.6;
ppCO2 (mmHg) — 4.1 (MCA freshly calibrated).
PMA-1: Shell heater temp (deg C) — 23.5
PMA-2: Shell heater temp (deg C) — 19.1
(n/a = data not available)

Propulsion System (PS): Total propellant load available (SM + FGB +
Progress) — 3689 kg (8133 lb) [as of 1/9/03].

Electrical Power Systems (EPS):
Both P6 channels fully operational. Beta Gimbal Assembly (BGA) 2B and
BGA 4B both in Autotrack (sun-following) mode.
SM batteries (as of 3am): Battery #7 is off line (failed); battery #4
is in "Cycle" mode; all other batteries (6) are in "Partial
Charge" mode.
FGB batteries (3am): Batteries #3 is offline (failed); all other batteries
(5) are in "Partial Charge" mode.
Plasma Contactor Unit PCU-1 and PCU-2 were active during the EVA. Back
on Standby afterwards.

Thermal Control Systems:
Air conditioner SKV-1 is Off, SKV-2 is Off.

Command & Data Handling Systems:
C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
GNC-1 MDM is prime; GNC-2 is Backup (new patches loaded on both).
EXT-2 is On (primary), EXT-1 is off.
LA-1, LA-2 and LA-3 MDMs are all operating.
PL-1 MDM is On (primary); PL-2 MDM is off (diagnostic
APS-1 (automated payload switch #1) and APS-2 are both On.
SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational. Lane
1 is down (as of 11/14).
SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.

Attitude Source:
3 CMGs on-line (CMG-1 failed).
State vector — U.S. SIGI-1
Attitude — U.S. SIGI-1
Angular rates — U.S. RGA-1

Flight Attitude:
LVLH TEA (local vertical/local horizontal = "earth-fixed":
z-axis in local vertical, x-axis in velocity vector [yaw: -10 deg, pitch:
-9.1 deg, roll: 0 deg]) until 6:15pm EST tonight, then XPOP.

Communications & Tracking Systems:
FGB MDM-1 is powered Off; FGB MDM-2 is operating.
All other Russian communications & tracking systems are nominal.
S-band is operating nominally.
Ku-band is operating nominally.
Audio subsystem operating nominally.
Video subsystem operating nominally.
MCOR (medium-rate communications outage recorder) is operating nominally.

SSRMS/Canadarm2 based at Lab PDGF with Keep Alive (KA) power on both
MBS: KA power on both strings. MT: at WS7, with KA power. POA: KA power
on both strings.
RWS (robotics workstations): Lab RWS is Off; Cupola RWS is Off.

ISS Orbit (as of this morning, 7:07am EST [= epoch]):
Mean altitude — 390.1 km
Apogee — 393.9 km
Perigee — 386.3 km
Period — 92.36 min.
Inclination (to Equator) — 51.63 deg
Eccentricity — 0.00056
Orbits per 24-hr. day — 15.59
Altitude loss — 180 m (mean) in last 24 hours
Revolutions since FGB/Zarya launch (Nov. ’98) — 23718

For more on ISS orbit and worldwide naked-eye visibility dates/times,

SpaceRef staff editor.